### v1

# FedEFC: Federated Learning Using Enhanced Forward Correction Against Noisy Labels

Abstract
Federated Learning (FL) is a powerful framework for privacy-preserving distributed learning. It enables multiple clients to collaboratively train a global model without sharing raw data. However, handling noisy labels in FL remains a major challenge due to heterogeneous data distributions and communication constraints, which can severely degrade model performance. To address this issue, we propose FedEFC, a novel method designed to tackle the impact of noisy labels in FL. FedEFC mitigates this issue through two key techniques: (1) prestopping, which prevents overfitting to mislabeled data by dynamically halting training at an optimal point, and (2) loss correction, which adjusts model updates to account for label noise. In particular, we develop an effective loss correction tailored to the unique challenges of FL, including data heterogeneity and decentralized training. Furthermore, we provide a theoretical analysis, leveraging the composite proper loss property, to demonstrate that the FL objective function under noisy label distributions can be aligned with the clean label distribution. Extensive experimental results validate the effectiveness of our approach, showing that it consistently outperforms existing FL techniques in mitigating the impact of noisy labels, particularly under heterogeneous data settings (e.g., achieving up to 41.64% relative performance improvement over the existing loss correction method).
1 Introduction
Federated Learning (FL) is a powerful paradigm for distributed learning that enables the training of a high-performing global model without requiring the aggregation or centralization of locally stored data [13]. While FL provides strong privacy guarantees by keeping client data decentralized, this non-centralized nature makes the model highly sensitive to the underlying data distribution among clients. In particular, the challenge of convergence under heterogeneous data distributions has been extensively studied [11, 10, 6, 1].
Recently, a growing body of research [23, 22, 3, 24] has investigated the impact of noisy datasets in addition to data heterogeneity. The adverse effects of label noise—whether caused by natural annotation errors or adversarial attacks—are often more pronounced in FL than in centralized learning (CL) due to the decentralized nature of training and the aggregation of corrupted model updates. Moreover, the lack of direct access to client datasets significantly limits the applicability of conventional noise mitigation techniques commonly used in CL, necessitating novel approaches tailored to the FL setting.
Building on this perspective, we propose FedEFC, an effective FL algorithm for mitigating the impact of noisy datasets. Our approach utilizes two key techniques:
  • Prestopping : A dynamic early stopping mechanism that prevents overfitting to mislabeled data by halting training at an optimal point.
  • Loss Correction : A robust adjustment of model updates to account for label noise, ensuring mitigation of noisy labels after the prestopping point.
Here, the proposed loss correction method is applied after the prestopping phase, replacing the standard update. Notably, our loss correction technique is specifically designed for consistent effectiveness in heterogeneous FL settings by refining and extending the forward correction method proposed in [16]. To achieve this, we introduce an alternative estimation process that improves the accuracy of the noise transition matrix and dynamically updates the loss function, leading to an enhanced forward correction mechanism. The overall architecture of FedEFC is depicted in Fig. 1. Our main contributions in this work are as follows:
  • We develop an enhanced forward correction to mitigate the impact of noisy labels without directly altering the data, thereby preventing unnecessary information loss (Sec. 3). When integrated with the prestopping technique, this approach effectively reduces the adverse effects of noisy labels, particularly in heterogeneous FL settings.
  • We provide a theoretical proof demonstrating that the enhanced forward correction enables each client to achieve the comparable training effectiveness as if learning from entirely clean data, despite the presence of noisy labels. (Sec. 3).
  • Experimental results confirm that FedEFC outperforms existing FL techniques, demonstrating its robustness in mitigating the impact of noisy labels and ensuring reliable model performance (Sec. 4). To achieve this, we introduce sparsity, a measure that quantifies the degree of asymmetric label noise, into our noisy label synthesis process [15]. Additionally, we leverage the Dirichlet distribution and Bernoulli distribution to systematically allocate data in a heterogeneous manner, ensuring a realistic simulation of non-IID conditions in FL. (Sec. 2).
1.1 Related Works
Confident learning : As data-centric AI has gained prominence over model-centric approaches, effectively handling noisy labels has become increasingly critical, particularly when working with large-scale datasets [18, 12]. Numerous studies have investigated techniques for identifying and managing mislabeled data [21, 2, 4, 15]. Among these, confident learning [15] leverages a count matrix to model the relationship between true and noisy labels. This matrix has been demonstrated to be highly effective in refining mislabeled instances within noisy datasets. In this work, we incorporate the count matrix into FedEFC to further enhance the forward correction, improving robustness against label noise in heterogeneous setting.
In [24], count matrix has also been integrated into FL for label correction. However, this approach has notable limitations, as it does not explicitly account for heterogeneous data distributions across clients and depends on a pretrained model for reliable performance, which may not always be feasible in real-world FL scenarios. Though, since adapting confident learning in FL offers a meaningful baseline for comparison with FedEFC, we modify it to ensure a fair and consistent evaluation within FL setting.
Forward loss correction : forward correction [16] is one of the main approaches designed to mitigate the detrimental effects of noisy labels by adjusting model predictions based on an estimated noise transition matrix. Both theoretical analysis and experimental results have validated the effectiveness of loss correction, demonstrating that it enables training on noisy datasets to approximate the learning dynamics of training on clean datasets. Nevertheless, a major challenge remains in constructing the reliable noise transition matrix, which is crucial for effective correction. In this work, FedEFC utilizes forward loss correction for the clients’ local training of FL, while the estimation of noise transition matrix is tailored to be robust in the heterogeneous FL settings. To enhance reliability, the matrix is re-estimated at each training round after the prestopping point, leveraging the temporal global model. Furthermore, for estimation, we apply the manner of count matrix, instead of the model prediction. As shown later, this method provides a more stable estimation, particularly in heterogeneous environments. A detailed discussion of this approach is provided in Sec. 3.
FL algorithms for non-IID settings : In real-world FL scenarios, data is typically distributed in a non-IID manner, introducing significant challenges for model training [5, 25]. Various strategies have been proposed to enhance FL performance under heterogeneous settings. For instance, FedProx [10] introduces minor modifications to FedAvg to achieve more robust convergence when training on non-IID data. FedDyn [1] dynamically aligns the local optimization objectives with the global loss function, ensuring more stable updates across clients. Ditto [11] enhances data robustness and fairness, improving individual client performance while maintaining overall model consistency. In this work, we consider these methods as baseline algorithms for comparison with FedEFC, evaluating its effectiveness in mitigating noisy labels and handling heterogeneous FL settings.
FL algorithms against noisy dataset : Several research efforts have explored ways to resolve the noise labels present in local datasets of FL. In [22], FedCorr is designed to cope with noisy data challenges. On the other hand, the framework requires the assumption that certain clients have entirely clean data to ensure improvement. RHFL [3] proposes a robust noise loss function for noisy labels under non-IID whereas public data is necessary to utilize the proposed algorithm. Although RoFL [23] effectively leverages similarity-based learning to mitigate the issue of asymmetric noisy labels, it requires additional information on feature data that may pose a potential risk to privacy and communication bottleneck. Of the approaches considered, FedCorr is compared with the proposed method as it serves as a more suitable baseline for a fair evaluation. Unlike methods relying on additional public data (RHFL) or feature-based similarity learning (RoFL), FedCorr better aligns with the constraints and challenges of real-world FL scenarios.
Figure 1: Overview of FedEFC framework. The scheme consists of two phases: (1) determining the prestopping point and (2) refining loss correction. In Phase 1, the centralized server tracks client training accuracies to identify the prestopping point where model parameters are near-optimal. In Phase 2, each client updates its loss function using enhanced forward correction, guiding global model parameters toward their optimal configuration in the clean data space.
2 Preliminaries
In this section, we describe the fundamentals of FL, including weight aggregation, data allocation, and noise generation, as applied in an FL scenario. In particular, we focus on non-IID data allocation and a practical noise generation method to reflect real-world settings.
In an FL environment, we assume a federated network with a centralized server and $N$ clients. The clients are elements of the set $\mathcal{S}$, where $|\mathcal{S}|=N$. Each client $k$ is assigned a local dataset denoted as $\mathcal{D}_{k}=\{(\bm{x}^{k},\tilde{y}^{k})\}^{n_{k}}$, consisting of $n_{k}$ examples. Each example pairs an input sample $\bm{x}^{k}$ with its observed noisy label $\tilde{y}^{k}$. The union of all local datasets is expressed as $\mathcal{D}=\bigcup_{k=1}^{N}\mathcal{D}_{k}$, but each client’s data remains private and is never shared with the server. All labels in the dataset belong to the label set $\mathcal{C}$, where $\tilde{y}^{k}\in\mathcal{C}$.
During each local training round, we define the local model weights for client $k$ as $w_{k}^{t}$, where $t$ represents the training round. The central server aggregates the weights from the participating clients in each round. The set of participating clients in round $t$ is denoted as $\mathcal{S}_{t}$ where $\mathcal{S}_{t}\subseteq\mathcal{S}$. Therefore, the global model weights $w_{g}^{t}$ at round $t$ are aggregated as follows, as described in [13]
$w_{g}^{t}\leftarrow\sum_{k\in\mathcal{S}_{t}}\frac{n_{k}}{\sum_{u\in\mathcal{S}_{t}}n_{u}}w_{k}^{t}.$ (1)
2.1 Data Allocation
We consider a non-IID data allocation strategy to reflect practical scenarios. Even in a non-IID setting, certain configurations can approximate IID characteristics by varying the balancing parameters. Two key parameters influence non-IID allocation: $\alpha_{\text{dir}}$ is derived from the Dirichlet distribution, which determines the number of examples assigned to each client, while $p$ is based on the Bernoulli distribution, which controls the label distribution among the examples allocated to clients. As described in [22], a combination of the Dirichlet and Bernoulli distributions is used to construct non-IID data distributions.
Specifically, the Bernoulli probability $p$ determines whether examples of a specific label $i$ are allocated to client $k$. The Bernoulli distribution is represented by the indicator $\mathbb{I}(k,i)\sim p$, which takes the value $1$ if label $i$ is allocated to client $k$, and 00 otherwise. Once the Bernoulli distribution is established, the number of examples assigned to clients with $\mathbb{I}(\cdot,i)=1$ follows a Dirichlet distribution parameterized by $\alpha_{\text{dir}}>0$. Therefore, the degree of non-IID data allocation is controlled by the parameters $p$ and $\alpha_{\text{dir}}$.
2.2 Noise Generation
The asymmetric noise flip more closely resembles real-world scenarios than the symmetric case when generating synthetic noise. To generate asymmetric noise, the sparsity mentioned in [15], [20] is utilized. The amount of noise $\rho$ and sparsity $\zeta$ need to be considered for noise generation. $\rho$ is the fraction of samples whose labels are flipped while $\zeta$ represents the proportion of uncorrupted labels, excluding the true label. For example, if all labels are flipped, $\zeta=0$. In contrast, $\zeta=1$ indicates that the dataset contains only clean labels. Thus, high sparsity represents greater imbalance in noisy labels.
3 Proposed Method
In this section, we propose FedEFC for mitigating the impact of noisy labels in non-IID FL. FedEFC consists of two phases: (1) determining a prestopping point and (2) applying loss correction. In Phase 1, the centralized server estimates the accuracy of the global model by aggregating the training accuracies of local clients. During Phase 1, each client transmits its measured accuracy to the centralized server, which monitors accuracy variations to determine the prestopping point. Once this point is reached, the server notifies the clients that Phase 1 is complete and that no further accuracy updates will be transmitted in subsequent rounds. In Phase 2, participating clients generate a noise transition matrix by analyzing the allocated dataset and apply loss correction. Detailed descriptions of each phase are provided below.
3.1 Finding the Prestopping Point
In Phase 1, we explore the learning property when the dataset includes noisy labels. As observed in [19], the loss of a model tends to sharply increase after a certain epoch when trained on a dataset with noisy labels. This phenomenon indicates that beyond prestopping point, the model can no longer effectively learn from clean data and instead begins to overfit to noisy labels [9]. As the accuracy of the global model fluctuates and saturates, clients experience a degraded learning performance on their local datasets. However, heuristic validation from a client’s perspective is difficult to adopt in FL. Therefore, we introduce the heuristic validation method suitable for an FL system.
We denote the estimated accuracy of the global model in round $t$ as follows
$A(t)=\sum_{k\in\mathcal{S}_{t}}\frac{A_{k}(w_{g}^{t})}{|\mathcal{S}_{t}|},$ (2)
where $A_{k}(w_{g}^{t})$ is training accuracy of client $k$ with the global model based on its own dataset. The training set contains both clean and noisy labels, which differs from the original heuristic validation approach that relies on a clean set [19]. A threshold $\gamma_{\text{thr}}$ is defined to quantify the instability in learning. In addition, a patience parameter $\tau_{p}$ is employed to track previous accuracy history. The accuracy measured in the current round, $A(t)$, is compared to that of the previous round, $A(t-1)$. If $A(t)<A(t-1)$, the patience parameter $\tau_{p}$ is incremented by 1; otherwise, $\tau_{p}$ is reset to 00. When the accuracy does not improve for $\gamma_{\text{thr}}$ consecutive rounds (i.e., $\tau_{p}=\gamma_{\text{thr}}$), the current round is regarded as the prestopping point, denoted as $T_{e}$. At the end of Phase 1, the centralized server instructs the clients to apply loss correction for every round after $T_{e}$. Empirical verification of the estimated accuracy, aggregated from each client’s training performance, is provided in the left panel of Fig. 2. In this panel, the estimated accuracy exhibits a sudden decline, attributed to the presence of noisy labels.
Figure 2: Left: Test accuracy and estimated accuracy used to determine the prestopping point. Right: Cosine similarity between the real noise matrix and the noise matrices estimated by the pretrained model and the global model in training. Experiments are conducted on the CIFAR-10 with the following settings: $\alpha_{\text{dir}}=10.0$, $p=0.5$, $\rho=0.2$, $\zeta=0.8$, and $T_{e}=54$.
3.2 Applying Loss Correction
In Phase 2, clients involved in rounds after $T_{e}$ apply loss correction using enhanced forward correction. Before applying loss correction, clients generate a noise transition matrix based on their own datasets. Each element of a noise transition matrix represents the probability of the true label given the observed label. The columns of this matrix correspond to the true labels, while the rows represent the observed labels. True labels are inferred with high confidence by the learning model whereas observed labels, which are potential noisy labels, are annotated in the dataset beforehand. Each client applies loss correction using the noise transition matrix. This loss correction allows the loss function to operate as if the dataset were noise-free.
3.2.1 Generating the Noise Transition Matrix
The clients receiving the Phase 1 completion signal generate the noise transition matrix. The noise transition matrix is derived from the count matrix, which consists of the number of examples for true and observed labels, as proposed in [15]. The key difference between these matrices lies in whether they represent conditional probability or label quantities. The count matrix is constructed with high reliability in a client-wise manner. The main challenge of FedEFC is that it requires a pretrained model. Without a pretrained model, the count matrix generated by the model currently being trained may consist of improper elements.
Figure 3: Example of generating the count matrix $C_{\tilde{y}|y}$. The figure illustrates the process for three labels—cat, dog, and hen—where “labeled” indicates data annotated with the observed label, and the true label is determined based on the threshold $\tau_{\text{label}}$.
However, in the FL setting, utilizing the global model trained up to the prestopping point helps improve the alignment between the estimated and real noise matrices, as demonstrated in the right panel of Fig. 2. At the prestopping point, the cosine similarity confirms that the noise matrices predicted by the pretrained model and the FL global model are nearly identical to the real noise matrix. Thus, the noise transition matrix is constructed by applying the global model after the prestopping point.
Each client acquires the noise transition matrix as follows. For client $k$, the global model after round $T_{e}$ predicts the true label as
$j=\arg\max_{\tilde{y}\in\mathcal{C}}\hat{p}(\tilde{y};\bm{x}^{k},w^{t}_{g}),$ (3)
where $\hat{p}_{k}(\cdot)$ denotes the estimated probability for client $k$. We define the set of examples with true label $y=j$ and observed label $\tilde{y}=i$ as
$\mathbf{X}_{\tilde{y}^{k}=i,y^{k}=j}$
=$\quad\quad=\Bigl{\{}\bm{x}^{k}\in\mathbf{X}_{\tilde{y}^{k}=i}:\hat{p}_{k}(\tilde{y}^{k}=j;\bm{x}^{k},w^{t}_{g})\geq\tau_{j}\Bigr{\}},$ (4)
where the set $\mathbf{X}_{\tilde{y}^{k}=i}$ represents all data samples with observed label $i$ in the dataset $\mathcal{D}_{k}$ of client $k$. The example set $\mathbf{X}_{\tilde{y}^{k}=i}$ satisfies
$\mathbf{X}_{\tilde{y}^{k}=i,y^{k}=j}\subseteq\mathbf{X}_{\tilde{y}^{k}=i}\subseteq\mathcal{D}_{k}.$ (5)
The threshold $\tau_{j}$ serves as the criterion for counting examples and is defined as
$\tau_{j}=\frac{1}{\left|\mathbf{X}_{\tilde{y}^{k}=j}\right|}\sum_{\bm{x}^{k}\in\mathbf{X}_{\tilde{y}^{k}=j}}\hat{p}_{k}(\tilde{y}=j;\bm{x}^{k},w^{t}_{g}),$ (6)
Threshold $\tau_{j}$ is the mean estimation probabilities for data labeled as $j$. Using the example set $\mathbf{X}_{\tilde{y}^{k}=i,y^{k}=j}$, the count matrix for client $k$ is defined as
$\bm{C}_{\tilde{y}^{k}=i,y^{k}=j}=\left|\mathbf{X}_{\tilde{y}^{k}=i,y^{k}=j}\right|.$ (7)
For simplicity, the notation is rewritten as
$\bm{C}_{\tilde{y}^{k}=i,y^{k}=j}\coloneqq\bm{C}^{k}_{\tilde{y}=i,y=j},\mathbf{X}_{\tilde{y}^{k}=i,y^{k}=j}\coloneqq\mathbf{X}^{k}_{\tilde{y}=i,y=j}.$ (8)
The count matrix is constructed to enumerate the number of examples associated with each label pair, with columns representing true labels and rows indicating observed labels. Fig. 3 illustrates an example of count matrix construction. The noise transition matrix of client $k$ is computed as
$\bm{Q}^{k}_{\tilde{y}=i|y=j}=\frac{\bm{C}^{k}_{\tilde{y}=i,y=j}}{\sum_{j\in\mathcal{C}}\bm{C}^{k}_{\tilde{y}=i,y=j}}.$ (9)
$\bm{Q}^{k}_{\tilde{y}=i|y=j}$ represents the $(i,j)$th element of the noise transition matrix $\bm{Q}^{k}_{{\tilde{y}}|{y}}$. This element corresponds to the conditional probability $p(\tilde{y}=i\mid y=j)$, which indicates the probability that the observed noisy label $\tilde{y}$ is $i$ given that the true label $y$ is $j$. Consequently, each client generates its own noise transition matrix $\bm{Q}^{k}_{{\tilde{y}}|{y}}$.
3.2.2 Applying Loss Correction
Participating clients apply loss correction based on the derived matrix $\bm{Q}_{{\tilde{y}}|{y}}$, following a similar approach to [16], where loss functions are modified. Regarding client $k$ in round $t$, the loss function $\mathcal{L}_{k}(\cdot)$ is given by
$\mathcal{L}_{k}(\bm{x},y;{w}_{k}^{t}).$ (10)
When a client trains with a DNN and applies the softmax function $\sigma_{k}(\cdot)$ in the final layer, the probability of the observed label $\tilde{y}=i$ given input $\bm{x}^{k}$ is
$\hat{p}_{k}(\tilde{y}=i|\bm{x}^{k};w_{k}^{t})$ $={\sigma_{k}}(\bm{h}_{i}^{k}(\bm{x}^{k};w_{k}^{t}))$ (11)
$=\frac{{\exp}({\bm{h}_{i}^{k}(\bm{x};w_{k}^{t})})}{\sum_{j\in\mathcal{C}}\exp({\bm{h}_{j}^{k}(\bm{x};w_{k}^{t})})}.$ (12)
$\bm{h}_{i}^{k}$ denotes the output of the DNN at the final layer corresponding to label $i$. The loss function $\mathcal{L}_{k}$ can be rewritten using cross-entropy as
$\mathcal{L}_{k}(\hat{p}_{k}(\tilde{y}^{k}=i|\bm{x}^{k};w_{k}^{t}))=-\log\hat{p}_{k}(\tilde{y}=i|\bm{x};w_{k}^{t}).$ (13)
To streamline notation, we replace $\bm{x}^{k}$ and $y^{k}$ with $\bm{x}$ and $y$, respectively. Using the noise transition matrix $Q_{\tilde{y}=i|{y}=j}^{k}$, the loss function $\mathcal{L}_{k}^{\textbf{F}}(\cdot)$ is updated as follows:
$\mathcal{L}_{k}^{\textbf{F}}(\hat{p}_{k}(\tilde{y}=i|\bm{x};w_{k}^{t}))=-\log{\hat{p}_{k}(\tilde{y}=i|\bm{x};w_{k}^{t})}$
$=-\log\left\{\sum_{j\in\mathcal{C}}p(\tilde{y}={i}|{y}=j)\cdot\hat{p}_{k}({y}=j|\bm{x};w_{k}^{t})\right\}$ (14)
$=-\log\left\{\sum_{j\in\mathcal{C}}Q^{k}_{\tilde{y}=i|y=j}\cdot\hat{p}_{k}({y}=j|\bm{x};w_{k}^{t})\right\}$ (15)
$=-\log\left\{Q_{\tilde{y}=i|{y}}^{k}\cdot\hat{p}_{k}({y}|\bm{x};w_{k}^{t})\right\}.$ (16)
$Q_{\tilde{y}=i|{y}}^{k}$ denotes the row vector of the noise transition matrix $Q^{k}_{{\tilde{y}}|{y}}$. The enhanced forward correction is obtained by weighting the loss function with the noise transition matrix $Q_{\tilde{y}=i|{y}}^{k}$. Clients then begin transmitting their learning parameters, updated via loss correction, to the centralized server. The entire process is summarized in Algorithm 1.
Algorithm 1 FedEFC: Federated Learning Using Enhanced Forward Correction
1:Global model $w_{g}^{0}$, maximum global epochs $T$
2:Final global model $w_{g}^{T}$
3:Phase 1: Find prestopping point
4:for $t=0$ to $T_{e}$ do
5:     $\mathcal{S}_{t}=Sampling({1,2,...,N})$; $\tau_{p}=0$
6:     for each client $k\in\mathcal{S}_{t}$ in parallel do
7:         $\left(w_{k}^{t+1},A_{k}(w_{g}^{t})\right)$leftarrow Client update($k$, $w_{g}^{t}$)
8:         Compute $A(t)$ through Eq. (2)
9:         if $A(t)>A({t-1})$ then
10:              $\tau_{p}\leftarrow 0$
11:         else
12:              $\tau_{p}\leftarrow\tau_{p}+1$
13:              if $\tau_{p}==\gamma_{\text{thr}}$ then
14:                  return $T_{e}=t$
15:              end if
16:         end if
17:     end for
18:     $w_{g}^{t+1}\leftarrow\sum_{k\in\mathcal{S}_{t}}\frac{n_{k}}{\sum_{u\in\mathcal{S}_{t}}n_{u}}w_{k}^{t+1}$
19:end for
20:Phase 2: Apply loss correction
21:for $t=T_{e}$ to $T$ do
22:     for each client $k\in\mathcal{S}_{t}$ in parallel do
23:         Compute ${Q}^{k}_{y=j|\tilde{y}=i}$ via Eq. (9)
24:         Update $\mathcal{L}_{k}^{\textbf{F}}(\bm{x},y;w_{k}^{t})$ using Eq. (3.2.2)
25:         $w_{k}^{t+1}$leftarrow Client update($k$, $w_{g}^{t}$, $\mathcal{L}_{k}^{\textbf{F}}(\cdot)$)
26:     end for
27:     $w_{g}^{t}\leftarrow\sum_{k\in\mathcal{S}_{t}}\frac{n_{k}}{\sum_{u\in\mathcal{S}_{t}}n_{u}}w_{k}^{t}$
28:end for
3.3 Objective Function Analysis in Noisy Labels
We analyze the objective function of FL in noisy labels when applying FedEFC. Each participating client updates its model by minimizing the local loss function during each round. Clients optimize their models iteratively. When using stochastic gradient descent (SGD), the update rule is given by ${w}_{k}^{t+1}\leftarrow{w}_{k}^{t}-\eta_{k}\nabla\mathcal{L}_{k}(\bm{x},y;{w}_{k}^{t})$ where $\eta_{k}$ denotes the learning rate of client $k$. According to [13], the objective function of FL is formulated as
$\arg\min_{w}$ ${F}_{g}({w})=\sum_{k\in\mathcal{S}_{t}}\frac{n_{k}}{\sum_{u\in\mathcal{S}_{t}}n_{u}}\mathcal{L}_{k}(\bm{x},y;w).$ (17)
The objective function in FL with noisy labels can be formulated using a composite loss, which combines a proper loss with a link function. A proper loss is used for class probability estimation, while a link function maps the classifier’s output to the range $[0,1]$. A composite loss is a proper loss, as proven in [17]. In the FL system, local loss function corresponds to a proper loss and the inverse of the softmax function acts as the link function ($\sigma^{-1}:[0,1]\rightarrow\mathbb{R}$). Since composite loss is defined as the combination of a proper loss and a link function, the composite loss for client $k$ is expressed as
$(\mathcal{L}_{k})^{\sigma^{-1}_{k}}({y},{\bm{h}^{k}(\bm{x}};w))=\mathcal{L}_{k}({y},\sigma_{k}({\bm{h}^{k}(\bm{x};w}))).$ (18)
In the case of a clean dataset, the minimizer of the loss function, following the property of composite proper losses, is given by
$\arg\min_{\bm{h}}\mathbb{E}_{\bm{x},{y}}[(\mathcal{L}_{k})^{\sigma^{-1}_{k}}({y},\bm{h}(\bm{x};w))]$ (19)
$=\sigma_{k}^{-1}(p({y}|\bm{x};w)).$ (20)
Therefore, the FL objective function can be rewritten as
$\arg\min_{w}{F}_{g}({w})$
$=\sum_{k\in\mathcal{S}_{t}}\frac{n_{k}\arg\min_{\bm{h}}\mathbb{E}_{\bm{x},{y}}[(\mathcal{L}_{k})^{\sigma^{-1}_{k}}({y},\bm{h}(\bm{x};w))]}{\sum_{u\in\mathcal{S}_{t}}n_{u}}.$ (21)
For a noisy dataset, the minimizer of the loss function is replaced by $\sigma_{k}^{-1}(p(\tilde{y}|\bm{x};w))$. However, under FedEFC, the objective function evaluated at the minimizer for clean data remains identical to that evaluated at the minimizer for noisy data when using enhanced forward correction. Theorem 1 demonstrates this equivalence, as derived in [16].
Theorem 1.
Let every $Q_{{\tilde{y}}|{{y}}}^{k}$ generated by client $k$ be a non-singular matrix. A composite loss incorporating $Q^{k}_{{\tilde{y}}|{y}}$ is given by
$(\mathcal{L}_{k}^{\textbf{F}})^{\sigma^{-1}_{k}}({y},{\bm{h}^{k}(\bm{x}};w))$ $=\mathcal{L}_{k}({y},Q^{k}_{{\tilde{y}}|{y}}\cdot\sigma_{k}({\bm{h}^{k}(\bm{x};w})))$
Then, the objective function evaluated at the minimizer for clean data is equivalent to that evaluated at the minimizer for noisy data:
$\sum_{k\in\mathcal{S}_{t}}\frac{n_{k}\arg\min_{\bm{h}}\mathbb{E}_{\bm{x},{\tilde{y}}}[(\mathcal{L}_{k}^{\textbf{F}})^{\sigma^{-1}_{k}}({y},\bm{h}(\bm{x};w))]}{\sum_{u\in\mathcal{S}_{t}}n_{u}}$
$=\sum_{k\in\mathcal{S}_{t}}\frac{n_{k}\arg\min_{\bm{h}}\mathbb{E}_{\bm{x},{y}}[(\mathcal{L}_{k})^{\sigma^{-1}_{k}}({y},\bm{h}(\bm{x};w))]}{\sum_{u\in\mathcal{S}_{t}}n_{u}}$
Proof. : We analyze the objective function by examining the minimizer of the loss function for client $k$. The minimizer of the loss function using the enhanced forward correction is derived as follows using $\phi_{k}=Q^{k}_{{\tilde{y}}|{y}}\cdot\sigma_{k}$:
$\arg\min_{\bm{h}}\mathbb{E}_{\bm{x},{\tilde{y}}}[(\mathcal{L}_{k}^{\textbf{F}})^{\sigma^{-1}_{k}}({y},\bm{h}(\bm{x};w))]$
$=\arg\min_{\bm{h}}\mathbb{E}_{\bm{x},{\tilde{y}}}[(\mathcal{L}_{k})^{\phi^{-1}_{k}}({y},\bm{h}(\bm{x};w))]$ (22)
$=\phi^{-1}_{k}(p({\tilde{y}}|\bm{x}))$ (23)
$=\sigma_{k}^{-1}\left((Q^{k}_{{\tilde{y}}|{y}})^{-1}\cdot p({\tilde{y}}|\bm{x})\right)$ (24)
$=\sigma_{k}^{-1}\left((Q^{k}_{{\tilde{y}}|{y}})^{-1}\cdot Q^{k}_{{\tilde{y}}|{y}}\cdot p({{y}}|\bm{x})\right)$ (25)
$=\sigma_{k}^{-1}\left(I\cdot p({{y}}|\bm{x})\right)$ (26)
$=\arg\min_{\bm{h}}\mathbb{E}_{\bm{x},{y}}[(\mathcal{L}_{k})^{\sigma^{-1}_{k}}({y},\bm{h}(\bm{x};w))].$ (27)
For client $k$, the minimizer of the loss function using enhanced forward correction under noisy labels is shown to be equivalent to the minimizer of the loss function under clean labels. Since the summation of the minimizers forms the objective function, Theorem 1 holds.
4 Experiments
In this section, FedEFC is evaluated on non-IID data with noisy labels using the MNIST [8], CIFAR-10 [7], and CIFAR-100 [14] datasets. To assess robustness against noisy labels, noise parameters $\rho$ and $\zeta$ are varied under fixed heterogeneity. Additionally, experiments are conducted with various heterogeneous parameters $\alpha_{\text{dir}}$ and $p$ while keeping noise levels constant to simulate environments ranging from nearly IID to highly non-IID. By combining these parameters, diverse scenarios with varying $(\alpha_{\text{dir}},p)$ and $(\rho,\zeta)$ are constructed. We compare FedEFC with standard FL techniques, noise-robust FL approach, and CL for DNNs trained with noisy labels. The results demonstrate that FedEFC is robust against both noisy labels and data heterogeneity.
4.1 Experiment Setup
Non-IID with Noisy Labels : In a real-world scenario, noise is first introduced into the training dataset before it is allocated to all clients in the FL system. Since similar noisy labels occur across clients, a consistent noise pattern is applied before partitioning the dataset according to non-IID parameters $(\alpha_{\text{dir}},p)$. This procedure creates a non-IID environment in which noisy labels are class-dependent.
Baselines : The performance of FedEFC is compared with several FL techniques: FedAvg [13], FedDitto [11], FedDyn [1], and FedProx [10]. FedCorr [22], a noise-robust FL method designed to tackle noisy labels, is included. Furthermore, once the proposed prestopping point is determined, both forward correction [16] and confident learning [15] are applied in the FL setting for comparison. As a reference, FedAvg without noise is employed to validate the robustness of the noise by comparing performance against the ideal noise-free case.
4.2 Implementation Details
FL experiments are simulated with $N=100$ clients. Each client performs local training using SGD with a momentum of $0.5$ and at each round, a fraction $0.1$ of clients are randomly selected for aggregation. The dataset-specific parameters are summarized in Table 1. The prestopping threshold, $\gamma_{\text{thr}}$, is set to $3$ for MNIST and $6$ for CIFAR-10 and CIFAR-100. To ensure training stability, prestopping monitoring begins after round $10$ for MNIST and after round $40$ for CIFAR-10 and CIFAR-100.
In cases of severe label imbalance, particularly when $p=0.2$, estimating the noise transition matrix directly from the raw count matrix is challenging. To compensate for this issue, FedEFC refines the count matrix by weighting it with the class distribution of each client before normalizing its columns. This adjustment is applied to MNIST and CIFAR-10 when $p=0.2$, while the standard count matrix remains effective under other settings.
Table 1: Dataset and FL parameter settings.
Parameter MNIST CIFAR-10 CIFAR-100
Class Size 10 10 100
Number of Clients 100 100 100
Model 9-Layers CNN ResNet-18 ResNet-34
Rounds 100 200 300
Local Batch Size 64 10 10
Table 2: Average test accuracy and standard deviation over 3 trials under noise-fixed settings with noise amount $\rho=0.2$ and sparsity $\zeta=0.8$ for various non-IID parameter configurations. The best two values in each column are highlighted in bold, except for the noiseless case (FedAvg wo noise).
Dataset MNIST CIFAR-10 CIFAR-100
($\alpha_{\text{dir}}$, $p$) (100.0, 0.8) (10.0, 0.5) (1.0, 0.2) (100.0, 0.8) (10.0, 0.5) (1.0, 0.2) (100.0, 0.8) (10.0, 0.5) (1.0, 0.2)
FedAvg $97.07\pm 0.10$ $96.31\pm 0.42$ $84.13\pm 2.23$ $82.57\pm 0.55$ $78.49\pm 0.75$ $61.27\pm 2.81$ $55.30\pm 0.23$ $54.97\pm 0.31$ $46.09\pm 0.35$
FedDitto $97.00\pm 0.05$ $96.44\pm 0.39$ $85.26\pm 3.61$ $83.53\pm 0.15$ $80.48\pm 0.95$ $59.53\pm 3.78$ $53.59\pm 0.37$ $53.91\pm 0.26$ $45.27\pm 0.15$
FedDyn $97.14\pm 0.08$ $96.11\pm 0.57$ $84.68\pm 1.83$ $82.64\pm 0.15$ $78.45\pm 0.84$ $60.25\pm 3.94$ $55.21\pm 0.38$ $54.39\pm 0.75$ $45.91\pm 0.64$
FedProx $97.11\pm 0.12$ ${96.25\pm 0.58}$ $84.76\pm 2.34$ $82.34\pm 0.57$ $78.77\pm 0.71$ $61.50\pm 2.16$ $55.68\pm 0.71$ $54.24\pm 0.43$ $44.73\pm 0.69$
FedCorr $96.47\pm 0.64$ $93.61\pm 1.16$ $90.69\pm 0.73$ ${84.90\pm 0.08}$ $72.54\pm 4.86$ $63.71\pm 0.73$ $\bm{60.11\pm 0.98}$ $\bm{57.31\pm 0.31}$ $44.80\pm 0.42$
confident learning $\bm{98.97\pm 0.13}$ $\bm{98.94\pm 0.16}$ $91.44\pm 2.75$ $\bm{88.01\pm 0.24}$ $80.21\pm 3.29$ $62.45\pm 1.21$ ${57.92\pm 0.11}$ $56.68\pm 0.83$ $47.49\pm 1.08$
forward correction ${97.05\pm 0.52}$ $94.79\pm 0.09$ $\bm{97.00\pm 0.20}$ $82.76\pm 0.21$ $\bm{80.62\pm 0.70}$ $\bm{70.80\pm 1.59}$  $56.72\pm 0.58$ $56.73\pm 0.22$ $\bm{52.00\pm 0.94}$
FedEFC $\bm{98.15\pm 0.45}$ $\bm{98.51\pm 0.46}$ $\bm{96.78\pm 0.60}$ $\bm{88.37\pm 0.08}$ $\bm{86.78\pm 0.36}$ $\bm{70.57\pm 1.25}$ $\bm{60.57\pm 0.40}$ $\bm{59.34\pm 0.51}$ $\bm{50.58\pm 0.18}$
FedAvg wo noise $99.49\pm 0.01$ $99.51\pm 0.02$ $98.91\pm 0.21$ $90.74\pm 0.14$ $88.93\pm 0.46$ $80.43\pm 0.86$ $68.14\pm 0.51$ $67.42\pm 0.32$ $63.58\pm 0.20$
Table 3: Average test accuracy and standard deviation over 3 trials under non-IID fixed settings with Dirichlet $\alpha_{\text{dir}}=10.0$ and Bernoulli $p=0.5$ for various noise configurations.
Dataset MNIST CIFAR-10 CIFAR-100
($\rho$, $\zeta$) (0.4, 0.8) (0.2, 0.4) (0.1, 0.0) (0.4, 0.8) (0.2, 0.4) (0.1, 0.0) (0.4, 0.8) (0.2, 0.4) (0.1, 0.0)
FedAvg $66.63\pm 0.26$ $96.94\pm 0.12$ $98.74\pm 0.02$ $58.42\pm 0.24$ $79.22\pm 0.29$ $84.83\pm 0.21$ $38.04\pm 0.59$ $55.45\pm 1.06$ $64.55\pm 0.49$
FedDitto $66.70\pm 0.31$ $97.09\pm 0.16$ $98.64\pm 0.04$ $59.18\pm 0.44$ $80.11\pm 0.72$ $83.09\pm 0.09$ $38.02\pm 0.40$ $54.26\pm 0.61$ $61.95\pm 0.47$
FedDyn $67.30\pm 0.33$ $97.04\pm 0.13$ $98.78\pm 0.02$ $58.36\pm 0.52$ $78.34\pm 0.69$ $84.75\pm 0.14$ $38.33\pm 0.64$ $55.51\pm 0.42$ $64.20\pm 0.29$
FedProx $67.13\pm 0.31$ $97.03\pm 0.12$ $98.82\pm 0.09$ $58.78\pm 0.14$ $79.24\pm 0.31$ $85.03\pm 0.06$ $38.18\pm 0.40$ $55.68\pm 0.67$ $64.82\pm 0.22$
FedCorr $65.71\pm 0.49$ $97.36\pm 0.60$ $\bm{99.19\pm 0.13}$ $54.06\pm 27.01$ $78.10\pm 1.42$ $80.31\pm 2.65$ $\bm{44.86\pm 0.07}$ $\bm{58.13\pm 1.57}$ $64.50\pm 1.78$
confident learning $\bm{68.38\pm 1.75}$ $\bm{99.22\pm 0.04}$ $\bm{99.33\pm 0.01}$ $\bm{66.20\pm 1.40}$ $\bm{84.95\pm 0.22}$ $\bm{86.83\pm 0.26}$ $41.37\pm 1.26$ $56.19\pm 0.62$ $64.51\pm 0.24$
forward correction ${67.69\pm 0.84}$ ${97.45\pm 0.28}$ $98.84\pm 0.08$ ${60.61\pm 1.47}$ ${80.76\pm 0.59}$ ${85.38\pm 0.11}$ $41.30\pm 0.27$ $57.59\pm 0.51$ $\bm{65.02\pm 0.07}$
FedEFC $\bm{71.35\pm 3.68}$ $\bm{98.39\pm 0.33}$ ${99.04\pm 0.04}$ $\bm{85.85\pm 0.66}$ $\bm{85.79\pm 0.09}$ $\bm{87.12\pm 0.16}$ $\bm{47.17\pm 0.59}$ $\bm{59.58\pm 0.41}$ $\bm{65.80\pm 0.33}$
FedAvg wo noise ${99.51\pm 0.02}$ ${88.93\pm 0.46}$ ${67.42\pm 0.32}$
4.3 Comparison with Baselines
Non-IID variation : To evaluate the impact of non-IID conditions, we fix noise at $(\rho,\zeta)=(0.2,0.8)$ and vary $(\alpha_{\text{dir}},p)$ across three configurations: $(100.0,0.8)$, $(10.0,0.5)$, and $(1.0,0.2)$ (see Table 2). FedEFC achieves the best performance in most configurations. By contrast, FedCorr achieves peak accuracy on MNIST and CIFAR-10 under nearly IID conditions $(100.0,0.8)$ but still lags behind FedEFC and suffers a more significant drop as non-IID characteristics intensify. forward correction performs optimally on MNIST, CIFAR-10, and CIFAR-100 in the most non-IID scenario $(1.0,0.2)$, as it leverages prediction probabilities for all labels, even in their absence from local datasets. However, its accuracy remains closely aligned with FedEFC, differing by at most 0.3% on MNIST and CIFAR-10 and under 1.5% on CIFAR-100. Similarly, confident learning achieves the best accuracy on MNIST under nearly IID $(100.0,0.8)$ and moderately non-IID $(10.0,0.5)$ conditions, yet its performance gap with FedEFC remains within 1%. Despite these cases, FedEFC maintains overall superiority, delivering more stable and resilient performance in non-IID FL environments.
Noise variation : Noise robustness is examined by fixing data allocation and varying noise settings across $(\rho,\zeta)$ pairs: high noise $(0.4,0.8)$, moderate noise $(0.2,0.4)$, and low noise $(0.1,0.0)$, as summarized in Table 3. In most noise conditions, FedEFC demonstrates greater robustness, consistently surpassing the baseline methods, with only minor differences in a few cases. For MNIST under low noise conditions, FedCorr and confident learning marginally outperform FedEFC, with differences of less than 0.3%. Similarly, for MNIST under moderate noise, confident learning attains the highest accuracy, but remains within a 1% margin of FedEFC. Beyond these minor variations, FedEFC consistently yields the best results. Notably, under high noise conditions on CIFAR-10, FedEFC experiences only a 3% performance drop compared to FedAvg in a noise-free setting.
5 Discussion
Enhanced forward correction : When applying the forward correction in FL, the key difference compared to FedEFC lies in how the noise transition matrix is generated. In FedEFC, the noise transition matrix is derived by counting data that satisfies a model confidence-based threshold, whereas forward correction computes the noise transition matrix directly from the predicted probabilities. In [16], the 97th percentile of the predicted probabilities is used to construct a concrete noise transition matrix. Deriving the noise transition matrix from predicted probabilities does not account for errors in the model predictions, whereas the count-based approach is more resilient to prediction errors, as shown in [15]. In FL with non-IID data, certain labels may be absent from local datasets, leading to reduced accuracy. To mitigate this weakness, class distribution weights are applied to the count matrix, as described in the Implementation Details.
The Nature of the global model in FL : A significant challenge in generating a noise transition matrix is the conventional requirement for a pretrained model. When a model in training is used to construct the noise transition matrix, it tends to overfit to the same dataset used for training the model. For instance, [2] employs an iterative noise cross-validation (INCV) approach that partitions the dataset into separate training and validation set to prevent improper noise transition matrix. In an FL system, the global model obtained prior to local client training is not overfitted to any particular client’s data. As a result, it can effectively generate the noise transition matrix without requiring a pretrained model or data partitioning, unlike conventional methods. FedEFC leverages the inherent properties of FL to overcome this challenge without additional procedures.
6 Conclusion
We propose FedEFC, a novel algorithm to address the noisy label problem in FL systems with non-IID data distributions. It consists of two phases: (1) finding a prestopping point and (2) applying loss correction. A key advantage of FedEFC is that it effectively mitigates the impact of noisy labels without requiring any inter-client information exchange. In particular, theoretical analysis establishes that, under FedEFC, the FL objective with noisy labels is equivalent to that with clean labels. Extensive experiments across varying noise levels and data heterogeneity demonstrates that FedEFC consistently outperforms conventional FL algorithms. Given that FedEFC operates under realistic conditions with noisy labels and non-IID data, it holds strong potential for adoption in diverse FL scenarios.